Calculate the number of liters of solvent needed to dissolve 44 grams of NaCl to create a 1.5 M solution
1
Expert's answer
2020-05-08T14:11:44-0400

molar mass of NaCl =58.5 g

moles of NaCl provided = 44/58.5 =0.75

"molarity= \frac{moles}{volume \ in \ litres }"


"1.5=\frac{0.75}{V}"

V=0.5 litres